# Coordinated Inauthentic Behavior on TikTok: Challenges and Opportunities for Detection in a Video-First Ecosystem

### v1

@@ Abstract
Detecting coordinated inauthentic behavior (CIB) is central to the study of online influence operations. However, most methods focus on text-centric platforms, leaving video-first ecosystems like TikTok largely unexplored. To address this gap, we develop and evaluate a computational framework for detecting CIB on TikTok, leveraging a network-based approach adapted to the platform’s unique content and interaction structures. Building on existing approaches, we construct user similarity networks based on shared behaviors, including synchronized posting, repeated use of similar captions, multimedia content reuse, and hashtag sequence overlap, and apply graph pruning techniques to identify dense networks of likely coordinated accounts. Analyzing a dataset of 793K TikTok videos related to the 2024 U.S. Presidential Election, we uncover a range of coordinated activities, from synchronized amplification of political narratives to semi-automated content replication using AI-generated voiceovers and split-screen video formats. Our findings show that while traditional coordination indicators generalize well to TikTok, other signals, such as those based on textual similarity of video transcripts or Duet and Stitch interactions, prove ineffective, highlighting the platform’s distinct content norms and interaction mechanics. This work provides the first empirical foundation for studying and detecting CIB on TikTok, paving the way for future research into influence operations in short-form video platforms.
@@ Introduction
With over 1.6B active users globally, [^1] https://www.statista.com/statistics/272014/global-social-networks-ranked-by-number-of-users/ TikTok has become one of the most influential social media platforms, playing a central role in shaping public discourse—especially among younger audiences. While its short-form video format offers an engaging space for creativity and self-expression, it has also become a vector for influence campaigns aiming to sway public opinion, disseminate disinformation, and manipulate political narratives. Such campaigns are particularly concerning during high-stakes events like national elections or public health crises, where coordinated dissemination of false or polarizing content can exacerbate societal divisions, trigger panic, or erode trust in institutions. Recent internal audits and journalistic investigations have highlighted TikTok’s growing role in covert influence operations, [^2] https://www.tiktok.com/transparency/en/covert-influence-operations?enter˙method=category˙card  [^3] https://www.cnn.com/2024/09/24/tech/tiktok-russian-media-accounts-us-election  [^4] https://www.nbcnews.com/tech/tech-news/tiktok-says-disrupted-15-influence-operations-year-one-china-rcna153831 underlining the urgent need for systematic study and effective monitoring of the platform’s information dynamics.
@@
Coordinated inauthentic behavior (CIB) is widely recognized as a hallmark tactic in online influence campaigns, where multiple accounts, often automated or inauthentic, amplify narratives in a synchronized fashion (Pacheco et al. 2021; Luceri et al. 2024; Tardelli et al. 2024a). These accounts often post similar content in close temporal proximity, reuse media assets, or repeatedly interact with one another to boost engagement metrics. While not inherently malicious, such coordination becomes problematic when it is deceptive in nature, e.g., when inauthentic accounts present themselves as genuine individuals or manipulate visibility through algorithmic gaming. These tactics are commonly used to promote specific political agendas, manufacture the illusion of consensus, or suppress opposing viewpoints, particularly during moments of societal stress such as elections, protests, or crises (Vishnuprasad et al. 2024; Alizadeh et al. 2020; Cinus et al. 2025).
Research on detecting CIB has largely focused on Twitter/$\mathbb{X}$ (Magelinski, Ng, and Carley 2022; Pacheco, Flammini, and Menczer 2020; Tardelli et al. 2024b; Luceri, Boniardi, and Ferrara 2024), where behavioral traces like co-retweets, shared URLs, and posting synchronicity have proven useful for uncovering coordinated efforts like state-linked information operations. A smaller body of work has investigated coordination on other platforms, such as Facebook and Telegram (Giglietto, Righetti, and Marino 2020; Giglietto et al. 2020; Cinus et al. 2025; Ng, Bartulovic, and Carley 2024). In contrast, video-first platforms like TikTok remain understudied, despite their growing relevance in the global information ecosystem.
To our knowledge, no study has yet systematically examined CIB on TikTok using computational methods. This gap is particularly important given the algorithmic design of TikTok’s For You feed, which personalizes content based on behavioral signals like viewing time, likes, and shares—rather than social connections or follower graphs. This design amplifies the reach of content regardless of follower count and makes it easier for coordinated campaigns to achieve high visibility if they trigger the right interaction patterns (Bandy and Diakopoulos 2020). As a result, a coordinated set of accounts can potentially manipulate visibility by posting similar content or interacting in patterns that exploit the algorithm’s engagement-based logic. Unlike Twitter/$\mathbb{X}$, where retweet networks and reply chains offer structural clues about coordination, TikTok lacks explicit interaction graphs, and reposting is relatively rare. Moreover, TikTok’s dominant use of video and audio complicates the direct application of text-based similarity methods that have proven effective on text-dominant platforms like Twitter/$\mathbb{X}$, Facebook, and Telegram. As a result, the detection of coordination on TikTok requires methodological adaptations tailored to its media format and engagement mechanics.
@@
To address this gap, we develop and evaluate a computational framework for detecting CIB on TikTok, leveraging a network-based approach adapted to the platform’s unique content and interaction structures. Building on established methodologies developed for Twitter/$\mathbb{X}$ (Luceri et al. 2024; Pacheco et al. 2021; Cinus et al. 2025), we first construct similarity networks that link users based on shared behaviors—such as uploading videos within tightly bounded time windows, or repeating similar captions and video content. Then, we apply graph pruning techniques to isolate dense subnetworks of potentially coordinated inauthentic accounts. Leveraging this approach, we analyze a dataset of 793K TikTok videos related to the 2024 U.S. Election (Pinto et al. 2025) to explore how inauthentic coordination manifests in the TikTok ecosystem.
Our findings reveal both promising opportunities and notable challenges in adapting CIB detection methods to TikTok. Several behavioral indicators previously validated on other platforms proved effective in uncovering clusters of likely coordinated activity on TikTok. These clusters exhibited patterns consistent with campaign-style amplification, including synchronized dissemination of political content, repeated posting of videos with visually identical formats and AI-generated voiceovers, and the use of split-screen formats to replicate messaging across accounts while circumventing content moderation. In many cases, accounts shared overlapping hashtag sequences and promoted the same external domains, including monetization platforms and mock websites—a commonly observed pattern on other platforms (Minici et al. 2024). [^5] https://www.veraai.eu/posts/uni-urbino-on-cib-around-pope-health-research-findings Additional signals of inauthenticity included auto-generated or extremely similar usernames, frequent handle alterations and swapping, and high-volume posting of near-duplicate content, suggesting semi-automated or orchestrated account control.
However, not all indicators yielded meaningful results. Traditional cues, such as textual similarity, and novel TikTok-specific features, such as reply, duet, and stitch, did not produce clusters indicative of suspicious coordination. These results likely reflect the platform’s unique norms of content creation and interaction, where reuse and remixing are common and often organically driven.
These observations underscore the challenges of applying coordination detection frameworks across platforms. While some TikTok-specific features did not yield strong evidence of coordination in our study, this does not imply their irrelevance. Instead, it points to the need for further methodological refinement and sustained evaluation of platform-specific signals. At the same time, our framework uncovered novel signatures of coordination unique to TikTok, such as watermark reuse, synthetic voice patterns, and the strategic use of manufactured split-screen videos. These findings suggest that TikTok-native signals hold significant potential for future detection efforts and should be incorporated into computational models designed to capture the evolving nature and increasingly sophisticated, AI-driven dynamics of inauthentic behaviors on video-first platforms.
Despite mounting evidence of coordinated manipulation, platform enforcement appears limited—only a small fraction of the suspicious accounts identified through our analysis were suspended or removed. This gap underscores the urgent need for proactive detection and moderation strategies. By systematically evaluating a diverse range of behavioral signals on a large-scale, election-related TikTok dataset, this study provides the first empirical foundation for detecting coordination on the platform and offers concrete directions for future research in this emerging area.
@@
Misinformation on TikTok, particularly in multimodal formats, presents additional challenges. Shang et al. (2021) address this issue through TikTec, a framework designed to detect multimodal COVID-19 misinformation in manipulated TikTok videos. Their findings emphasize the importance of integrating visual features, caption-guided attention, speech signals, and visual-speech co-attention to improve detection performance. Building on this work, Shang et al. (2025) introduce MultiTec, a system for identifying misinformation in short videos that leverages caption-guided visual representations, acoustic-aware speech modeling, and a dual-attentive transformer that combines visual-speech co-attention with multi-head sequential attention.
Along similar lines, Corso, Pierri, and Morales (2024) examined the prevalence and spread of conspiracy theories on TikTok. Their large-scale analysis identified a lower-bound prevalence of approximately 0.1% for conspiracy-laden videos, revealing how even a small fraction of such content can leverage TikTok’s algorithmic infrastructure to achieve wide reach. The study underscores the unique challenges posed by conspiratorial narratives in short-form, remixable formats, which often evade detection and contribute to the broader misinformation ecosystem.
Similarly, Zenone and Caulfield (2022) analyzed TikTok videos related to the 2022 monkeypox outbreak and identified videos propagating 11 types of conspiracy theories. Their findings highlight the importance of social media monitoring as a tool for detecting and addressing public health misinformation. Continued efforts to identify and mitigate misinformation on platforms with highly diverse, global audiences are becoming increasingly essential. Notably, TikTok creators themselves have reported frustration with the insufficient moderation of harmful comments and misleading content, raising further concerns about the platform’s role in content governance (Zeng and Kaye 2022).
@@
CIB has emerged as a key behavioral pattern underpinning the spread of disinformation, propaganda, and conspiratorial narratives across social media platforms (Pacheco, Flammini, and Menczer 2020; Luceri et al. 2024). While coordination is not inherently malicious, it is often observed in legitimate grassroots organizing and social movements, its deceptive forms are typically characterized by manipulative amplification of highly partisan or misleading content (Pacheco et al. 2021; Nizzoli et al. 2021).
@@
Among the most effective paradigms are network-based coordination detection techniques, which model user behaviors as similarity graphs constructed from shared hashtags, URLs, posting times, or interaction targets (Pacheco, Flammini, and Menczer 2020; Pacheco et al. 2021; Magelinski, Ng, and Carley 2022; Nizzoli et al. 2021). These similarity networks enable the application of topological measures, such as edge weights capturing behavioral similarity and centrality metrics to identify influential or densely connected nodes, to isolate clusters of users potentially involved in influence operations (Luceri et al. 2024; Minici et al. 2025; Mannocci et al. 2024; Tardelli et al. 2024b; Vishnuprasad et al. 2024). Much of this literature has focused on Twitter, leveraging its structured interaction data (e.g., retweets, replies, and quote tweets) to trace coordination across accounts. Parallel work has explored intra- and inter-platform CIB dynamics on different social media networks, such as Twitter, Facebook, Telegram, YouTube, and 4chan, also examining how content and actors migrate between ecosystems (Cinus et al. 2025; Ng, Cruickshank, and Carley 2022; Zannettou et al. 2019; Gatta et al. 2023).
In contrast, our work introduces a fine-grained, unsupervised CIB detection framework for TikTok, a platform that has thus far received little attention in this domain. By inspecting traditional and TikTok-specific behavioral indicators, this paper represents the first systematic attempt to extend network-based CIB detection to TikTok, providing both methodological innovations and empirical insights into coordinated influence on short-form video platforms.
@@
This study leverages the public TikTok 2024 U.S. Presidential Election Dataset released by Pinto et al. (2025), a large-scale, multimodal collection capturing TikTok video content and associated engagement metadata in the lead-up to the 2024 U.S. Presidential Election. The dataset was assembled using a combination of the TikTok Research API and third-party tools, with videos filtered through a curated and evolving set of election-related hashtags and keywords aligned with trending discourse throughout the collection period. The full dataset spans the period from November 1, 2023, to October 16, 2024, and contains over 3.1 million videos posted by more than 693K unique users.
For this study, we analyze a narrower subset of videos published between August 1 and September 30, 2024, a critical pre-election window marked by intensified campaign activity and heightened potential for coordinated messaging. This subset comprises approximately 793K videos posted by 238K unique users, capturing key events such as the official nomination of Kamala Harris (August 22) and the first presidential debate (September 10). In addition to video content and metadata, we extracted comments associated with these videos. For each video ID, comments were retrieved using a third-party crawler, resulting in structured data that includes the video ID, comment ID, comment text, and the username of the commenter. It is important to note that comment retrieval may fail for videos that were deleted, made private, or associated with accounts suspended at the time of extraction.
To support our coordination detection framework, we extract a diverse set of behavioral traces from this multimodal dataset, including video content, metadata, and user interactions, to construct user similarity networks. These behavioral signals form the basis of our detection strategy and are discussed in detail in the following section.
@@
Our coordination detection framework builds on network-based methodologies established in prior research (Pacheco et al. 2021; Luceri et al. 2024; Cinus et al. 2025), and is adapted to reflect the unique content modalities and interaction dynamics of the TikTok platform. The core assumption underlying this approach is that anomalously high behavioral similarity between users, across content, timing, or interaction patterns, can indicate underlying coordination, particularly when such overlap is unlikely to occur organically (e.g., due to exogenous factors, such as major external events, that could steer collective behavior).
Accordingly, our pipeline detects such patterns through three main stages: (1) the identification and extraction of behavioral traces that may indicate coordination; (2) the construction of user-entity bipartite graphs for each behavioral signal, followed by their projection into user-user similarity networks; and (3) the detection of CIB clusters through graph pruning techniques that apply thresholds on edge weights and node centrality scores.
@@
We define a set of behavioral traces that capture different dimensions of user similarity on TikTok. While some of these traces parallel those used on platforms like Twitter/$\mathbb{X}$, others are specifically adapted to account for TikTok’s platform-specific modalities and interaction norms. These behavioral signals, extracted from video content, descriptions, metadata, and comments, form the basis for constructing user-entity bipartite graphs across various coordination signals:
@@
  • Hashtag Sequence: Users posting videos that contain the same set of hashtags in the same exact order within their descriptions. Identical hashtag sequences, especially when repeated across multiple accounts, can signal coordinated attempts to influence discoverability and visibility in TikTok’s algorithmically curated feed.
@@
  • Synchronized Posting: Users who upload similar videos within narrow time windows, typically a few minutes apart, suggesting potential coordination. This behavior may indicate that multiple accounts are operating in a synchronized manner, through automation or centralized control, to maximize visibility and impact by flooding the platform with similar content at the same time.
  • Text Similarity: Users posting videos or comments with highly similar or identical content, which may suggest orchestrated messaging or script-based content generation. To enable analysis of spoken language in videos, we first transcribe the audio using OpenAI’s speech-to-text model, Whisper. High text similarity between users, particularly when occurring repeatedly, may indicate centrally coordinated messaging strategies.
@@
  • Co-Domain: Users who include links to the same external domains in their video descriptions or comments. Repeated linking to identical domains, particularly low-credibility, campaign-related, or monetized websites, can signal coordination, especially when combined with other behavioral similarities. To account for superficial variations (e.g., URL parameters), we extract and normalize the domain-level information (e.g., stripping query strings or subpaths), allowing us to detect shared linking behavior even when full URLs differ.
@@
  • Co-Duet and Co-Stitch: TikTok allows users to combine their videos with others using Duet or Stitch. In a Stitch, the video from another user is played first, followed by the user’s own content. In a Duet, the screen is split into two parts, showing the original and user-generated videos side by side. This feature is commonly used to engage with content from celebrities, media outlets, and viral creators. These interactions are typically indicated in the video description with phrases such as “Stitch with @username” or “Duet with @username.”
@@
  • Co-Reply: TikTok allows users to respond to comments through video replies. These replies display the original comment as a text overlay and are marked by descriptions like “Replying to @username.” While often used for organic engagement, this feature can potentially be exploited for coordinated activity—e.g., by repeatedly replying to comments from associated accounts to boost content visibility and engagement.
@@
For each behavioral trace, we construct a bipartite graph where nodes represent users on one side and behavioral entities on the other (e.g., hashtag sequences, domains, or time bins). An edge exists between a user and an entity if the user engages with that entity. For instance, a user is connected to a specific hashtag sequence if they post a video with that exact ordered set of hashtags in the description, or a user is linked to a temporal bin (e.g., a 5-minute interval) if they post a video within that time window. For co-duet, co-stitch, and co-reply a user is linked to another user (as an entity) if they perform a duet, stitch, or reply interaction with their content.
@@
The Text Similarity trace follows a different procedure, as it does not involve bipartite graph construction. Instead, we compare textual content from video transcripts or comments by directly computing pairwise similarity scores. All textual inputs undergo a standardized cleaning process, including the removal of punctuation, stopwords, emojis, and URLs. We retain only text segments with a minimum length of four words, as shorter inputs tend to introduce noise and yield unreliable similarity scores. To compute embeddings, we use the multilingual sentence-transformer model stsb-xlm-r-multilingual from Hugging Face. Cosine similarity is then calculated using the FAISS library for scalable approximate nearest-neighbor search—performed on an RTX A6000 GPU. To optimize for computational efficiency and temporal relevance, we evaluate similarity only within a sliding one-day window and apply a similarity threshold of 0.7 to retain the most semantically aligned text pairs, in line with prior research (Vishnuprasad et al. 2024; Pacheco, Flammini, and Menczer 2020). The resulting similarity network connects users with edge weights corresponding to the average similarity score across all matching text pairs. This enables the detection of clusters of users disseminating near-duplicate or semantically equivalent content.
@@
Following pruning, we analyze the resulting network structures to assess whether clusters exhibit strong signals of CIB. We adopt a mixed-methods approach, combining automated analysis with manual inspection, and examine a wide range of features, e.g., repeated usernames, shared bios and profile images, or links pointing to the same suspicious domains. Importantly, the co-occurrence of multiple coordination signals increases confidence that a given cluster constitutes an orchestrated operation.

### v2

@@ Abstract
Detecting coordinated inauthentic behavior (CIB) is central to the study of online influence operations. However, most methods focus on text-centric platforms, leaving video-first ecosystems like TikTok largely unexplored. To address this gap, we develop and evaluate a computational framework for detecting CIB on TikTok, leveraging a network-based approach adapted to the platform’s unique content and interaction structures. Building on existing approaches, we construct user similarity networks based on shared behaviors, including synchronized posting, repeated use of similar speech segments, and multimedia content reuse, and apply graph pruning techniques to identify dense networks of likely coordinated accounts. Analyzing a dataset of 1.35M TikTok videos related to the 2024 U.S. Presidential Election, we uncover a range of coordinated activities, from synchronized amplification of political narratives and semi-automated content replication to AI-generated voiceovers and manufactured split-screen video formats. Our findings demonstrate that while several indicators proposed in this study effectively detect CIB on TikTok, other platform-native signals, such as video-based replies and Duet and Stitch interactions, often reflect organic engagement rather than inauthentic behavior, highlighting the platform’s distinct content dynamics and interaction mechanics. This work provides the first empirical foundation for studying CIB on TikTok, paving the way for future research into influence operations on video platforms.
@@ Introduction
With over 1.6B active users globally, [^1] https://www.statista.com/statistics/272014/global-social-networks-ranked-by-number-of-users/ TikTok has become one of the most influential social media platforms, playing a central role in shaping public discourse, especially among younger audiences. While its short-form video format offers an engaging space for creativity and self-expression, it has also become a vector for influence campaigns aiming to sway public opinion, disseminate disinformation, and manipulate political narratives. Such campaigns are particularly concerning during high-stakes events like national elections or public health crises, where coordinated dissemination of false or polarizing content can exacerbate societal divisions, trigger panic, or erode trust in institutions. Recent internal audits and journalistic investigations have highlighted TikTok’s growing role in covert influence operations, [^2] https://www.tiktok.com/transparency/en/covert-influence-operations?enter˙method=category˙card [^3] https://www.cnn.com/2024/09/24/tech/tiktok-russian-media-accounts-us-election [^4] https://www.nbcnews.com/tech/tech-news/tiktok-says-disrupted-15-influence-operations-year-one-china-rcna153831 underlining the urgent need for systematic study and effective monitoring of the platform’s information dynamics.
@@
Coordinated inauthentic behavior (CIB) is widely recognized as a hallmark tactic in online influence campaigns, where multiple accounts, often automated or inauthentic, amplify narratives in a synchronized fashion (Pacheco et al. 2021; Luceri et al. 2024; Tardelli et al. 2024a). These accounts often post similar content in close temporal proximity, reuse media assets, or repeatedly interact with one another to boost engagement metrics. While not inherently malicious, such coordination becomes problematic when it is deceptive in nature, e.g., when inauthentic accounts present themselves as genuine individuals or manipulate visibility through algorithmic gaming. These tactics are commonly used to promote specific political agendas, manufacture the illusion of consensus, or suppress opposing viewpoints, particularly during moments of societal stress (Vishnuprasad et al. 2024; Alizadeh et al. 2020; Cinus et al. 2025).
[…]
@@
[…]
However, not all indicators yielded meaningful results. TikTok-specific features, such as video-based reply, duet, and stitch, did not produce clusters indicative of coordination. These results likely reflect the platform’s unique norms of content creation and interaction, where reuse and remixing are common and often organically driven. While these TikTok-specific features did not yield strong evidence of CIB, our framework uncovered novel signatures of coordination unique to TikTok, such as watermark reuse, synthetic voice patterns, including audio deep fake, and the strategic use of manufactured split-screen videos. These findings suggest that TikTok-native signals hold significant potential for future detection efforts and should be incorporated into computational models designed to capture the evolving nature and increasingly sophisticated, AI-driven dynamics of inauthentic behaviors on video-first platforms.
Despite mounting evidence of coordinated manipulation, platform enforcement appears limited: Only a small fraction of the suspicious accounts identified through our analysis were suspended or removed. This gap underscores the urgent need for proactive detection and moderation strategies. By systematically evaluating a diverse range of behavioral signals on a large-scale, election-related TikTok dataset, this study provides the first empirical foundation for detecting coordination on the platform and offers concrete directions for future research in this emerging area.
@@
Misinformation on TikTok, particularly in multimodal formats, presents additional challenges (Shang et al. 2025). Corso, Pierri, and Morales (2024) examined the spread of conspiracy theories on TikTok. Their large-scale analysis identified a lower-bound prevalence of approximately 0.1% for conspiracy-laden videos, revealing how even a small fraction of such content can leverage TikTok’s algorithmic infrastructure to achieve wide reach. The study underscores the unique challenges posed by conspiratorial narratives in short-form, remixable formats, which often evade detection and contribute to the broader misinformation ecosystem.
Similarly, Zenone and Caulfield (2022) analyzed TikTok videos related to the 2022 monkeypox outbreak and identified videos propagating several conspiracy theories. Their findings highlight the importance of social media monitoring as a tool for detecting and addressing public health misinformation. Notably, TikTok creators have reported frustration with the insufficient moderation of harmful comments and misleading content, raising further concerns about the platform’s role in content governance (Zeng and Kaye 2022). To address this issue, Shang et al. (2021) propose TikTec, a framework designed to detect COVID-19 misinformation on TikTok. Their findings emphasize the importance of integrating visual features, captions, and speech signals to improve detection performance.
@@
CIB has emerged as a key behavioral pattern underpinning the spread of disinformation, propaganda, and conspiratorial narratives across social media platforms. While coordination is not inherently malicious, its deceptive forms are typically characterized by manipulative amplification of highly partisan or misleading content (Nizzoli et al. 2021).
@@
Among the most effective paradigms are network-based coordination detection techniques, which model user behaviors as similarity graphs constructed from shared hashtags, URLs, posting times, or interaction targets (Pacheco, Flammini, and Menczer 2020; Pacheco et al. 2021; Magelinski, Ng, and Carley 2022; Nizzoli et al. 2021). These similarity networks enable the application of topological measures, such as edge weights capturing behavioral similarity and centrality metrics to identify influential or densely connected nodes, to isolate clusters of users potentially involved in influence operations (Luceri et al. 2024; Minici et al. 2025; Mannocci et al. 2024; Tardelli et al. 2024b; Vishnuprasad et al. 2024). Much of this literature has focused on Twitter, leveraging its structured interaction data (e.g., retweets) to surface coordination. Parallel work has explored intra- and inter-platform CIB dynamics on different social media networks, such as Twitter, Facebook, Telegram, YouTube, and 4chan, also examining how content and actors migrate between ecosystems (Cinus et al. 2025; Ng, Cruickshank, and Carley 2022; Zannettou et al. 2019; Gatta et al. 2023).
In contrast, our work introduces a fine-grained, unsupervised CIB detection framework for TikTok, a platform that has thus far received little attention in this domain. By inspecting traditional, novel, and TikTok-specific behavioral indicators, this paper represents the first systematic attempt to extend network-based CIB detection to TikTok, providing both methodological innovations and empirical insights into coordinated influence on short-form video platforms.
@@
This study leverages the public TikTok 2024 U.S. Presidential Election Dataset released by Pinto et al. (2025), a large-scale, multimodal collection capturing TikTok video content shared in the lead-up to the 2024 U.S. Presidential Election. The dataset was assembled using the TikTok Research API, with videos filtered through a curated and evolving set of election-related hashtags and keywords (listed in the Appendix) aligned with trending discourse throughout the collection period. The full dataset spans the period from November 1, 2023, to January 20, 2025.
For this study, we analyze a narrower subset of videos published between August 1 and October 31, 2024, a critical pre-election window marked by intensified campaign activity and heightened potential for coordinated messaging. This subset comprises approximately 1.35M videos posted by 362K unique users, capturing key events such as the official nomination of Kamala Harris and the first presidential debate. Additionally, we retrieved video files and associated comments using a third-party crawler. We were able to successfully download comments for approximately 92% of the videos; some retrievals failed due to videos being deleted, made private, or shared by suspended accounts.
@@
Our coordination detection framework builds on network-based methodologies established in prior research (Pacheco et al. 2021; Luceri et al. 2024; Cinus et al. 2025), and is adapted to reflect the unique content modalities and interaction dynamics of the TikTok platform.
The core assumption underlying this approach is that anomalously high behavioral similarity between users, across content, timing, or interaction patterns, can indicate underlying coordination, particularly when such overlap is unlikely to occur organically (e.g., due to exogenous factors, such as major external events, that could steer collective behavior). Accordingly, our pipeline detects such patterns through three main stages: (1) the extraction of behavioral traces that may indicate coordination; (2) the construction of similarity networks for each behavioral signal; and (3) the detection of CIB clusters through graph pruning techniques.
@@
[…]
@@
  • Hashtag Sequence: Users posting videos with the same set of hashtags in the same exact order, especially when repeated across multiple accounts, can signal coordinated attempts to influence discoverability and visibility in TikTok’s algorithmically curated feed.
@@
  • Synchronized Posting: Users who post content within narrow time windows, typically a few minutes apart, may indicate that multiple accounts are operating in a synchronized manner, through automation or centralized control, to maximize visibility and impact by flooding the platform with similar content at the same time.
@@
  • Co-Domain: Users who repeatedly link to the same external domains in video descriptions (or comments) may be exhibiting coordinated behavior, especially when those domains are affiliated with low-credibility, campaign-related, or monetized websites. To account for superficial variations (e.g., URL parameters), we extract and normalize the domain-level information (e.g., stripping query strings or subpaths), allowing us to detect shared linking behavior even when full URLs differ.
@@
  • Co-Duet and Co-Stitch: TikTok allows users to combine their videos with others using Duet or Stitch. In a Stitch, the video from another user is played first, followed by the user’s own content. In a Duet, the screen is split into two parts, showing the original and user-generated videos side by side. While this feature is commonly used to engage with content from celebrities, media outlets, and viral creators, it may be used for coordinated attacks.
@@
  • Co-Reply: TikTok allows users to respond to comments through video replies. These replies display the original comment as a text overlay and are marked by descriptions like “Replying to @username.” While often used for organic engagement, this feature can potentially be exploited for coordinated activity, e.g., by repeatedly replying to comments from associated accounts to boost content visibility and engagement.
  • Speech Similarity: Users posting videos with highly similar or identical speech segments may indicate orchestrated messaging or script-based content generation. To enable analysis of spoken language in videos, we first transcribe the audio using OpenAI’s speech-to-text model, Whisper. High speech similarity between users, particularly when occurring repeatedly, may indicate centrally coordinated messaging strategies.
  • Video Similarity: Users who post visually or structurally similar videos may be participating in orchestrated content dissemination. We leverage visual models to identify instances of near-duplicate or templated content. High video similarity across multiple accounts, especially when observed repeatedly and within narrow temporal windows, is a strong indicator of coordinated amplification. This is particularly relevant on TikTok, where platform-native video editing tools (e.g., templates, split-screens, and overlays) can facilitate subtle forms of content reuse and signal campaign orchestration.
@@
We construct distinct similarity networks, each corresponding to one of the seven behavioral traces examined in this study. For each behavioral trace, except Speech and Video Similarity, we construct a bipartite graph where nodes represent users on one side and behavioral entities on the other (e.g., hashtag sequences, domains, or time bins). An edge exists between a user and an entity if the user engages with that entity. For instance, a user is connected to a specific hashtag sequence if they post a video with that exact ordered set of hashtags in the description, or a user is linked to a temporal bin (e.g., a 5-minute interval) if they post a video within that time window. For co-duet, co-stitch, and co-reply a user is linked to another user (as an entity) if they perform a duet, stitch, or reply interaction with their content.
@@
Speech Similarity and Video Similarity traces depart from the bipartite graph approach. Instead, we compute direct pairwise user similarity based on content embeddings derived from audio and visual inputs.
For the Speech Similarity behavioral trace, we compare textual content from the audio transcripts extracted with Whisper. All textual inputs undergo a standardized cleaning process, including the removal of punctuation and stopwords. We retain only speech segments with a minimum length of four words, as shorter inputs tend to introduce noise and yield unreliable similarity scores. To compute embeddings, we use the multilingual sentence-transformer model stsb-xlm-r-multilingual from Hugging Face. Cosine similarity is then calculated using the FAISS library for scalable approximate nearest-neighbor search. To isolate only the most synchronized and potentially inauthentic behaviors, we focus on user pairs exhibiting perfect similarity (cosine similarity = 1) and simultaneous posting (0-second difference), indicating both semantic and temporal alignment. To further reduce the likelihood of coincidental similarity, we require that such exact matches occur at least twice between the same user pair, consistent with prior work (Vishnuprasad et al. 2024; Pacheco, Flammini, and Menczer 2020). This conservative approach emphasizes repeated, deliberate coordination rather than one-off coincidences. A link is established between two users if they satisfy this criterion, resulting in a similarity network where edges connect users who have posted highly similar content at the exact same time on multiple occasions.
For the Video Similarity behavioral trace, we employed the ViSiL similarity model (Kordopatis-Zilos et al. 2019) to generate embeddings for all videos in the dataset. These embeddings were indexed using FAISS to enable efficient pairwise similarity computation at scale. Following a procedure analogous to the Speech Similarity trace, we constructed a video similarity network, where edges connect users who repeatedly posted near-duplicate videos in a highly synchronized manner. Specifically, we retained only video pairs with a similarity score exceeding 0.9 and a posting time gap of exactly 0 seconds. An edge between two users was created only if they shared at least two such highly similar videos, ensuring that the detected patterns reflect repeated and temporally aligned video reuse.
@@
Following pruning, we analyze the resulting network structures to assess whether clusters exhibit strong signals of CIB. We adopt a mixed-methods approach, combining automated analysis with manual inspection, and examine a wide range of features, e.g., repeated usernames, shared bios and profile images, and the co-occurrence of multiple coordination signals, such as synchronized posting, content reuse across modalities (e.g., hashtags, audio, visuals), and templated formats (e.g., split-screen or AI-generated voiceovers). Importantly, the co-occurrence of these behavioral and identity-level cues increases confidence that a given cluster is not the result of organic alignment but rather reflects likely coordinated inauthentic behavior. Manual inspection was performed independently by two annotators, consistently yielding exceptionally high agreement (see the Appendix for additional details).